An 80% fruit juice concentrate is to be mixed with water to produce 10 gallons of 20%
fruit juice mixture. How much of each is needed? Is it even legal to call the resulting
concoction “juice”?

Answer by stanbon(75887)

An 80% fruit juice concentrate is to be mixed with water to produce 10 gallons of 20%
fruit juice mixture. How much of each is needed?
----------------------
Equation:
water + water = water
-----
0.20x + 1*(10-x) = 0.80*10
------
Multiply thru by 100 to get:
20x + 100*10 - 100x = 80*10
-80x = -10*10
x = 1.25 gallons (amt. of 20% juice needed)
================
10-x = 8.75 gallons (amt. water needed)
===================================
Cheers,
Stan H.